OpenStudy (anonymous):

The surface area of the top surface of the water in a circular swimming pool is about 206 square feet. Estimate the radius of the pool, to the nearest foot.

OpenStudy (anonymous):

Well, the area of a circle is A = (pi) * (radius)(radius) also known as radius squared. If we just fill in the equation with the given area of 206. Then, 206 = (pi)(radius^2) (radius^2) = 206/(pi) radius = sqrt(206/pi) ft NOTE: This is the exact answer, not simplified. That I'll leave for you to do.